# Fast In-Spectrum Graph Watermarks

### v1

@@ Appendix D Scalability reduction
TABLE IV: Worst-time complexities of FFG to watermark a graph $G$ with $N$ vertices with a key of length $m$ with and without the dimensionality reduction to $N_{0}$ vertices.
D-A Guideline’s complexities
In Section IV-C, we give practical guidelines for low distortion. We now studied this operation in detail, and its effect on the scalability of FFG.
As mentioned before, we set the key length $m$ of FFG as the key length used in [13]. In fact, $m$ can be set in worst-case complexity $\mathcal{O}(N_{0})$. Once the key length $m$ is set, it is always possible to find a value $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and for any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. $\sigma_{max}$ can be estimated with at most $\mathcal{O}(\log N_{0})$ operations. An automated way to set $\sigma$ resulting in a small ED guarantee, is a dichotomous search in $[\![1,\sigma_{max}]\!]$. The worst-case time complexity of such setting is testing $\mathcal{O}(\sigma_{max}\log\sigma_{max})$ values for $\sigma$ when each test is of complexity $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. To summarize, the worst-case time complexity for setting $\sigma$ knowing $m$ is $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. This is the same as the worst-case complexity of the embedding function, so the proposed guidelines do not affect the scalability of FFG.
Appendix E Additional experiments
E-A (Goal 5) Testing undetectability w. Graph Neural Networks
In addition to the properties stated in the paper and which follow the properties stated in the related papers, we add a Goal 5. We want to measure whether a graph neural network is able to distinguish a watermarked graph from an original graph. Indeed, if such a network existes, it could be maliciously used by an adversary in addition to the classical edge-flip attack.
@@ Appendix E Additional experiments
We generated $2,000$ independent Barabási-Albert graphs of $1,000$ vertices with an attachment parameter of $3$. Subsequently, we watermarked half of these graphs by generating $1,000$ independent keys with parameters $m=200$ and $\sigma=100$ (chosen as outlined in Section IV-C to achieve an ED between $1$ and $100$ edges of the graph, i.e. between $0.034$ and $3.4$% edges). Our dataset includes the $1,000$ non-watermarked graphs (of label 0) and the $1,000$ watermarked graphs (of label 1). Notably, we underline that the seed graphs for generating the watermarked graphs were excluded from the dataset, to avoid bias in the classification task.
We adapted the graph classification tutorial from the Python library called PyTorch Geometric [30] with our dataset. The task of the GNN is to distinguish graphs into two classes: those who have been watermarked and those who were not. The degrees of the vertices have been added as their features in this dataset. In other words, the degree sequence of each graph will serve as a basis for attempts in detecting a deviation from the benign sequences of non watermarked graphs. The training is set to $80$% of the dataset, leaving $20$% of graphs as a test set.
Our experiments show that with its default parameters or some classic variants of these (number of hidden channels, learning rate…), the GNN is not able to succeed in the task of discriminating between watermarked graphs and unwatermarked graphs. Indeed the accuracy on train and test datasets oscillates between 0.4775 and 0.5225 depending on the hyperparameters.
We conclude that this attack is ineffective against FFG. This implies that the watermarks inserted in FFG are small enough (ED between $0.034$ and $3.4$% edges) to be undetectable by such a GNN classification attack.
E-B Testing Laplacian eigenvalue’s stability
In our examples, we focus on Edit Distance (ED) due to its generality and widespread use in the state-of-the-art. We also consider degree centrality, as shown in Figure 4. The next section discusses the impact of FFG on the Laplacian eigenvalues of a graph, a metric that is sometimes used in graph studies.
@@ E-B Testing Laplacian eigenvalue’s stability
In conclusion, we suggest that future readers test their own metrics of interest before sharing graphs watermarked with Cox-G. Even if there is no theoretical guarantee on the preserving of this metric, the variation induced for the watermark is sometimes negligible for the desired use.

### v2

@@ VI Conclusion
TABLE IV: Worst-time complexities of F&F to watermark a graph $G$ with $N$ vertices with a key of length $m$ with and without the dimensionality reduction to $N_{0}$ vertices.
Guideline’s complexities
In Section IV-C, we give practical guidelines for low distortion. We now studied this operation in detail, and its effect on the scalability of F&F.
As mentioned before, we set the key length $m$ of F&F as the key length used in [13]. In fact, $m$ can be set in worst-case complexity $\mathcal{O}(N_{0})$. Once the key length $m$ is set, it is always possible to find a value $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and for any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. $\sigma_{max}$ can be estimated with at most $\mathcal{O}(\log N_{0})$ operations. An automated way to set $\sigma$ resulting in a small ED guarantee, is a dichotomous search in $[\![1,\sigma_{max}]\!]$. The worst-case time complexity of such setting is testing $\mathcal{O}(\sigma_{max}\log\sigma_{max})$ values for $\sigma$ when each test is of complexity $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. To summarize, the worst-case time complexity for setting $\sigma$ knowing $m$ is $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. This is the same as the worst-case complexity of the embedding function, so the proposed guidelines do not affect the scalability of F&F.
-E Additional experiments
(Goal 5) Testing undetectability w. Graph Neural Networks
In addition to the properties stated in the paper and which follow the properties stated in the related papers, we add a Goal 5. We want to measure whether a graph neural network is able to distinguish a watermarked graph from an original graph. Indeed, if such a network existes, it could be maliciously used by an adversary.
@@
We generated 2,0002,0002 , 000 independent Barabási-Albert graphs of 1,0001,0001 , 000 vertices with an attachment parameter of 333. Subsequently, we watermarked half of these graphs by generating 1,0001,0001 , 000 independent keys with parameters $m=200$ and $\sigma=100$ (chosen to achieve an ED between 111 and 100100100 edges of the graph, i.e. between 0.0340.0340.034 and 3.43.43.4% edges). Our dataset includes the 1,0001,0001 , 000 non-watermarked graphs (of label 0) and the 1,0001,0001 , 000 watermarked graphs (of label 1). Notably, we underline that the seed graphs for generating the watermarked graphs were excluded from the dataset, to avoid bias.
We adapted the graph classification tutorial from the Python library called PyTorch Geometric [30] with our dataset. The task of the GNN is to distinguish graphs into two classes: those who have been watermarked and those who were not. The degrees of the vertices have been added as their features in this dataset. In other words, the degree sequence of each graph will serve as a basis for attempts in detecting a deviation from the benign sequences of non watermarked graphs. The training is set to 808080% of the dataset, leaving 202020% as a test set.
Experiments show that with its default parameters or some classic variants of these (number of hidden channels, learning rate…), the GNN does not succeed in the task of discriminating between watermarked graphs and unwatermarked graphs. Indeed the accuracy on train and test datasets oscillates between 0.4775 and 0.5225 depending on hyperparameters.
We conclude that this attack is ineffective against F&F. This implies that the watermarks inserted in F&F are small enough (ED between 0.0340.0340.034 and 3.43.43.4% edges) to be undetectable by such a GNN classification attack.
Testing Laplacian eigenvalue’s stability
In our examples, we focus on Edit Distance (ED) due to its generality and widespread use in the state-of-the-art. We also consider degree centrality, as shown in Figure 4. The next section discusses the impact of F&F on the Laplacian eigenvalues of a graph, a metric that is sometimes used in graph studies.
@@
In conclusion, we suggest that future readers test their own metrics of interest before sharing graphs watermarked with F&F. Even if there is no theoretical guarantee on the preserving of this metric, the variation induced for the watermark is sometimes negligible for the desired use.